Calculate the variance for the following sample of n 5 scores. If it is a decimal number, leave two numbers after the decimal point and do not round Variance-
Answer #1

Ans.

X​​​​​​i (X​​​​​​i - X bar)2
2 1
3 0
1 4
5 4
4 1
15 10

Xbar = 15/5 = 3

Variable(X) = 10/5 = 2